Operation Summary for May 14, 2010 - May 17, 2010 |
The Integrated Luminosity for the period from 5/10/10 to 5/17/10 was 61.08 pb-1. Friday May 14 The day shift began with Operations monitoring store 7811 (40.33E30) and the Recycler stash (317.3E10), with the Antiproton Source (Pbar) stacking (15.9E10), with the Linac, Booster, and the Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, Operations received the operational signoffs at 10 AM for the next Meson MTest (T1005) experiment to begin taking beam on Monday. Operations terminated store 7811 at 11:30 AM. Operations put the TeV into HEP shot setup and established store 7812 at 1:15 PM with an average initial luminosity (AIL) of 372E30. Operations reported a problem with a Booster RF station (BRF12); RF personnel adjusted the bias power supply ground fault circuit. A Switchyard person established beam to MTest at 2:31 PM. During the evening shift, Operations disabled the MTest beam at 6:02 PM. Operations reported a short amount of downtime due to problems with a Booster ion pump (B:IPRF17) and RF station (BRF15), and with a Linac RF station (LRF5). Saturday May 15 The midnight shift began with Operations monitoring store 7812 (97.0E30) and the stash (300.0E10), with Pbar stacking (14.0E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, at 12:30 AM, store 7812 aborted, which induced magnet quenches at sectors F1, D1, B0, and D0. Quench investigation began, but did not reveal a hardware related failure. Meanwhile, the D1 cryo system struggled to recover from the quench. Operators consulted CHL personnel about the vacuum in the D15-16 regions. Operations prepared the TeV A-E enclosures and Transfer Hall for accesses to address possible vacuum issues at D1. During the day shift, Mechanical Support personnel made an access into the TeV and reported a frosted and open Kautzky valve at D15-3. Cryo technicians had the valve repaired by noon and put the cryo system into recovery. Operations established beam to MTest at 2:34 PM. During the evening shift, the D1 cryo system recovered at 5:30 PM. During the TeV checkout and turn on, Operations reported that a D1 beam valve (D1BVU) would not open. Operations called in Mechanical Support personnel to investigate. Sunday May 16 The midnight shift began with TeV in access, with Operations monitoring the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations reported that efforts to open the upstream beam valve at D1 continued. Operators swapped valve controller cards without success. Operations called in Controls personnel to investigate. Efforts to exercise the valve in the tunnel proved difficult, as the valve was located behind the steel shielding inside the D0 collision hall (D1 side). They replaced a faulty solenoid and verified the remote control of the valve before reinstalling the detector shielding. Mechanical Support personnel completed their accesses by 4:30 AM. Operators began the TeV checkout and turn on. Operations had put the TeV through a dry squeeze and checked orbits by 6:15 AM. Operations put the TeV into shot setup at 6:30 AM and established store 7815 at 7:30 AM with an AIL of 384.8E30. Operations contacted Linac personnel about repeated quadrupole (Q5 Bulk) power supply trips. The Pbar Accumulator injection septa developed an internal power supply problem; operators contacted Pbar personnel to investigate. Operations had put the Pbar Source into standby during shot setup due to the increasing number of septa power supply trips. During the day shift, Pbar personnel replaced/repaired a bad 5V power supply, internal to the Accumulator injection septa, and resumed stacking at 9:30 AM. Operations established beam to MTest at 10:02 AM. Linac personnel arrived to investigate problems with the Q5 BULK power supply, but did not find any obvious hardware related problems. The power supply settled down following a double hardware swap with the Q4 Bulk power supply. The NuMI Lambertson (LAM61) tripped the NuMI beam permit via the C204 module several times. During the evening shift, Operations terminated store 7815 at 7:10 PM. Operations disabled the MTest beam at 6 PM. Operations put the TeV into shot setup and established store 7816 at 8:30 PM with an AIL of 341.3E30. The initial luminosity was lower than expected due to larger than normal Pbar longitudinal emittances. Monday May 17 The midnight shift began with Operations monitoring store 7816 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations reported that LAM61 continued to pull the NuMI beam permit due to a noisy C204 signal. Operations established beam to MTest at 7:09 AM. The Plans for Monday The plans for today are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 05/17/2010 email Fermilab |
| Security, Privacy, Legal | |